OpenStudy (anonymous):

Determine whether the equation has no, one, two or infinite solutions without graphing. y=x^2 y=x-1 y=(x-3)^2-1 y=x^2-6x+8

OpenStudy (anonymous):

 if you want to sure that y=x^2 and y=x-1 have not solution you must solve x^2=x-1 (equivalent to equations)